40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Mon Oct 3 13:59:40 2016, awade, DailyProgress, TempCtrl, Modified LIGO temperature sensor interface board 20161002_D980401_PSLTempBox.epstemp_AD590TransImpedanceBox_PSDs.zipD980400-B.pdf
    Reply  Mon Oct 3 18:15:23 2016, not awade, DailyProgress, TempCtrl, Modified LIGO temperature sensor interface board 
       Reply  Tue Oct 4 11:49:53 2016, awade, DailyProgress, TempCtrl, Modified LIGO temperature sensor interface board 20161002_D980401_PSLTempBox_Vrms.eps20161002_D980401_PSLTempBox_Vrms.fig
          Reply  Sun Oct 9 09:01:27 2016, not awade, DailyProgress, TempCtrl, Modified LIGO temperature sensor interface board 
             Reply  Mon Oct 10 11:32:27 2016, awade, DailyProgress, TempCtrl, Modified LIGO temperature sensor interface board 20161002_D980401_PSLTempBox.pdf
    Reply  Fri Sep 1 13:36:20 2017, awade, DailyProgress, TempCtrl, Modified LIGO temperature sensor interface board 
       Reply  Sun Sep 3 14:10:38 2017, awade, Kira, DailyProgress, TempCtrl, Modified LIGO temperature sensor interface board LIGO-Temperature-Sense-Board-SingleChannel.pdf
          Reply  Mon Sep 4 23:07:41 2017, Kevin, DailyProgress, TempCtrl, Modified LIGO temperature sensor interface board TemperatureSensingCircuit.pdfTemperatureBox.jpg
             Reply  Tue Sep 5 23:54:45 2017, Kevin, DailyProgress, TempCtrl, Modified LIGO temperature sensor interface board p1.jpgp2.JPG
          Reply  Wed Sep 6 19:58:21 2017, awade, Kira, DailyProgress, TempCtrl, Modified LIGO temperature sensor interface board 
Message ID: 1895     Entry time: Mon Sep 4 23:07:41 2017     In reply to: 1893     Reply to this: 1897
Author: Kevin 
Type: DailyProgress 
Category: TempCtrl 
Subject: Modified LIGO temperature sensor interface board 

I looked at the temperature sensing board that Andrew and Kira were working on (PSL:1889) that had ~12 kHz ringing when the op-amp was driving a 100 Ω load or smaller. I saw the same ringing except that it was ~830 kHz.

Since this ringing is likely caused by the load drawing too much current from the op-amp, Andrew suggested removing the second op-amp stage and adding a resistor in series with the output of the first op-amp; schematic for a single channel attached. After adding a 1 kΩ resistor, the ringing stopped.

The board Andrew and Kira were working on was having intermittent issues with the +12 V regulator which they had replaced; it looks like one of the pads could have been damaged. I modified one of the other temperature sensing boards to use this new circuit for all four sensing channels.

Attachment 1: TemperatureSensingCircuit.pdf  15 kB  | Hide | Hide all
TemperatureSensingCircuit.pdf
Attachment 2: TemperatureBox.jpg  248 kB  | Hide | Hide all
TemperatureBox.jpg
ELOG V3.1.3-